### v2

# Cargo Sherlock: An SMT-Based Checker for Software Trust Costs

Abstract.
Supply chain attacks threaten open-source software ecosystems. This paper proposes a framework for quantifying trust in third-party software dependencies that is formally checkable – formalized in satisfiability modulo theories (SMT) – while at the same time incorporating human factors, like the number of downloads, authors, and other metadata that are commonly used to identify trustworthy software in practice. We use data from software analysis tools and metadata to build a first-order relational model of software dependencies; to obtain an overall “trust cost” combining these factors, we propose a formalization based on the minimum trust problem, which asks for the minimum cost of a set of assumptions that can be used to conclude that the code is safe. We implement these ideas in Cargo Sherlock, targeted for Rust libraries (crates), incorporating a list of candidate assumptions motivated by quantifiable trust metrics identified in prior work. Our evaluation shows that Cargo Sherlock can identify synthetically generated supply chain attacks and known incidents involving typosquatted and poorly AI-maintained crates. Its performance scales to Rust crates with many dependencies.
Formal Methods, Trust in Software, Human Factors in Security, Software Supply Chain
journalyear: 2026copyright: ccconference: IEEE/ACM 14th International Conference on Formal Methods in Software Engineering; April 12–13, 2026; Rio de Janeiro, Brazilbooktitle: IEEE/ACM 14th International Conference on Formal Methods in Software Engineering (FormaliSE ’26), April 12–13, 2026, Rio de Janeiro, Brazildoi: 10.1145/3793656.3793687isbn: 979-8-4007-2478-7/2026/04ccs: Software and its engineering Formal methodsccs: Security and privacy Software and application securityccs: Security and privacy Human and societal aspects of security and privacy
1. Introduction
As software ecosystems grow in size and complexity, automated techniques are needed to prevent supply chain attacks (Ohm et al., 2020; Martínez and Durán, 2021). Recent and notable examples of such attacks include the SolarWinds attack (Martínez and Durán, 2021); the XZ Utils backdoor (Freund, 2024; Zorz, 2024); and for Rust, which is the focus of our implementation, the malicious libraries rustdecimal (Rust Secure Code Working Group, 2024b) and faster_log (Pearce, 2025). These attacks occur at the package granularity, where untrustworthy libraries or software dependencies are added to trustworthy projects, or safe libraries are hijacked by malicious developers. Many existing techniques can be used to help identify such attacks. Some of these are formally checkable in the sense that they are based on a well-defined formal model of software: e.g., static analysis tools, fuzzers, type systems, and verification tools. For such tools, analysis results correspond to concrete syntactic or semantic source code properties or program traces, and can therefore be understood in an appropriate formal system (Møller and Schwartzbach, 2012). However, these tools are often limited to specific semantic or syntactic properties of programs.
On the other hand, industry developers and security engineers typically rely on a rather different set of criteria in practice: they use a combination of trust in software and auditing to ensure that third-party dependencies are safe from software supply chain attacks (Hamer et al., 2025; Mozilla, 2024; Zoghbi et al., 2026; Chin, 2023; development team, 2024b). For example, developers at a large company may be likely to trust software that was developed by that company or one of its industry partners; at a smaller scale, an independent developer may be more inclined to trust software that is widely accepted by the community (e.g., has a large number of downloads and stars on GitHub). Trust and auditing are the foundation of techniques like software bill of materials (SBOM) (Muirí, 2019; Hendrick and Zemlin, 2022; Xia et al., 2023; Zahan et al., 2023), which require human-curated or manually generated lists of software components and their purposes. However, trust and auditing (as currently practiced) often come with no formal guarantees; there is no formal model over which we can understand the meaning of audits and trust relationships, and how they compose.
Research goals.
Is it possible to quantify trust in a way that is formally checkable (in some formal system) yet also accounts for human factors that many developers rely on in practice? Our specific focus is to quantify trust in propositional and first-order logic; roughly speaking, the more assumptions that must be taken for a software dependency “on faith”, the greater the trust cost that we should need to pay in order to believe it functions correctly. It should be the case that higher trust costs correspond to more risky code. We identify the following specific goals for such a trust cost:
  • Trust should be specific: a trust cost should commit the analysis to a specific assumption or set of assumptions about the software (e.g., the number of downloads, trusted audits, or the results of specific analysis tools).
  • Trust should be auditable: it should be possible for a human to read and understand the assumptions made by a trust analysis.
  • Trust should be configurable: the user should be able to add and revise assumptions, e.g., by adding additional tool outputs or human factors as parameters.
  • Finally, trust should be composable: we should be able to compare and combine trust costs corresponding to different portions of the package ecosystem.
We note that traditional software vulnerability analysis and other machine learning (ML) approaches (Krsul, 1998; Ghaffarian and Shahriari, 2017) are a non-solution to the above. Such ML techniques would meet our criteria for configurability and composability; but they would fail the first and second criteria, auditability and specific assumptions that can later be vetted by human developers and human users.
Cargo Sherlock: an SMT-based approach.
To address these goals, we propose Cargo Sherlock, a design and implementation for quantifying supply chain trust based on a formalization in satisfiability modulo theories (SMT). Cargo Sherlock can be used directly by end users to quantify trust when selecting dependencies to add to a project, and its features are motivated by this intended use case; see Figure 1 and Section 2.
We begin with an SMT-based formalization of trust that we call the minimum trust problem (Section 3). This particular formalization assumes that we have a list of assumptions that the user might make about the code (such as: “if the code has a large number of downloads, then it is safe” or “if the code is audited by a trusted organization, then it is safe”); each assumption has an associated cost, which can be thought of as the cost inherent in making that assumption. These assumptions are motivated by prior empirical research; importantly, they are not necessarily true facts (if they were true unconditionally then trust would not be required!), but rather, they are real thought processes that users might potentially employ to conclude that dependencies are safe, and are configurable by users. The minimum trust problem then is to determine the minimum cost of a set of assumptions, encoded in propositional logic, that logically entail that the code is safe. We argue that this model can meet the goals we outlined above; for example, it produces trust costs that are specific to a particular set of assumptions, can be configured to meet specific users’ needs, and are auditable by human developers. We consider two algorithms for solving the minimum trust problem, based on a reduction to SMT. We also show that the problem is NP-hard in the worst case. This formalization is similar in principle to existing work from the logic programming and databases communities, for example, the semiring framework for database provenance (Green and Tannen, 2017), assigning weights to logical conclusions in Datalog (Bistarelli et al., 2008), and solvers for weighted Max-SAT (Heras et al., 2008; Li and Manya, 2009).
For the design and implementation (Section 4), Cargo Sherlock gathers software supply chain metadata and the results of automated program analysis tools to build a first-order relational structure of the software supply chain – that is, of all the dependencies of a given package. It then instantiates the minimum trust problem above on a set of possible assumptions on the variables involved in this first-order structure. The implementation also supports “parameterized assumptions”, where the cost depends in a continuous way on a particular piece of metadata, and “negative assumptions”, which can be used to incorporate information about software components that are distrusted, such as known vulnerabilities or bugs found by the program analysis tools.
Our evaluation (Section 5) asks four questions: Can Cargo Sherlock be used to prevent synthetically generated supply chain attacks? Would Cargo Sherlock identify risky Rust dependencies (crates) that have been involved in known supply chain incidents? How does Cargo Sherlock fare on known vulnerable crates? Lastly, does the performance of Cargo Sherlock scale to Rust crates with many dependencies? Among our results, we find that synthetically generated typosquatting attacks generally result in much higher trust costs. We also find that of the two algorithms we propose, our more optimized algorithm provides timely results for crates with a large number of dependencies.
One way to put these results into context is to think about a particular user model of how we expect users to interact with our tool. For example, due to the higher trust costs, if using our tool in the prescribed model, a user would catch any of the typosquatting attacks considered in our evaluation by identifying and omitting the risky dependency prior to its inclusion.
Contributions in summary:
  • A formal model of the supply chain trust problem and two algorithms for its solution using the SMT theory of linear arithmetic (Section 3);
  • A design and architecture for a usable tool based on these ideas, including concrete input data about Rust code and a collection of built-in candidate assumptions motivated by human factors and available program analyses (Section 4);
  • An evaluation of the efficacy of our tool for identifying synthetic and real-world supply chain attacks, and its scalability on Rust libraries with many dependencies (Section 5).
We conclude the paper with related work (Section 6). Cargo Sherlock is open source (MIT license) and available via a virtual machine image on Figshare [^1] https://doi.org/10.6084/m9.figshare.31094197 and as a repository on GitHub. [^2] https://github.com/davispl/cargo-sherlock
2. Motivating Overview
Figure 1. Cargo Sherlock example output on the anyhow crate. (Note that lower trust cost, higher distrust cost is better.)
Example output of Cargo Sherlock on the anyhow-1.0.97 crate, showing that it is marked as SAFE due to presence of an audit.
This section presents a motivating overview of Cargo Sherlock from a user-facing perspective; the reader who is more interested in formal details may skip to Section 3. We consider a hypothetical example of Ferris, a Rust developer who wishes to install a Rust crate called aho_corasick (implementing the Aho-Corasick algorithm), and explore how Cargo Sherlock’s approach might help Ferris attain the goals laid out in the introduction.
2.1. Automating trust costs
After finding the crate aho_corasick from a Google search, Ferris faces a choice: do they trust the crate blindly, or do they first try to determine if the crate is safe to install on their system? If Ferris is a typical developer, they might employ a common middle ground by relying on various heuristics to argue that the crate is likely safe: e.g., that the crate has many downloads, or that it is developed by someone who is trustworthy. At its core, our tool supports Ferris by attempting to automate this reasoning. The premises of Ferris’ arguments are based on assumptions about the supply chain ecosystem (e.g., “the more downloads a crate has, the safer it is”); Cargo Sherlock includes a default list of such assumptions motivated by prior research (see Section 4.3). Ferris may find some assumptions more reasonable than others; we mark each assumption with a cost to reflect its reasonableness. If Ferris runs Cargo Sherlock on aho_corasick, it will present them with aho_corasick’s trust cost: the total cost of the argument whose premises have the lowest total cost to assume (informally, the strongest argument). After using Cargo Sherlock, Ferris determines that aho_corasick is marked as trusted, and that it was marked SAFE due to having a high number of downloads. See Figure 1 for an example output; the trust cost (lower is better) is shown at the top in green.
2.2. Trust should be specific
Now suppose that Ferris becomes concerned about trusting crates with a high number of downloads. To address this concern, Ferris hires an expert auditor, Crabby, who combs through every line of source code of aho_corasick. Since Ferris trusts Crabby’s expertise, Ferris assigns a very low cost to the assumption that “if the crate is audited by Crabby, it is safe” – lower than any other available assumption. Next, suppose Crabby comes back after the audit and reports that the crate is completely safe to use. In this scenario, it would seem that the previous assumptions used to mark the crate trusted – such as its high number of downloads or the author of the crate – are no longer relevant to Ferris. Since Cargo Sherlock only reports the minimum trust cost for a crate with the associated assumptions, it does not present other (potentially unhelpful) assumptions. In other words, Ferris is only presented with a single specific reason to trust the crate. In Figure 1, the specific assumptions are shown under “Assumptions for Trusting” in green.
2.3. Trust should be auditable and configurable
Unfortunately, Ferris later finds out that Crabby is untrustworthy. Since this renders the existing audit suspect, Ferris revisits the previous assumptions, and increases the trust cost for “if the crate is audited by Crabby, it is safe”. In turn, Cargo Sherlock will update its result: the trust cost for aho_corasick will increase, and a new minimum set of assumptions will be identified. This same mechanism can be used if Ferris wants to add new information to the tool (e.g., adding additional trusted authors and auditors).
2.4. Trust should be composable
During this process, Ferris discovers that aho_corasick depends on a second crate, memchr, that was previously overlooked. Since this could present an additional attack surface, Ferris believes that the trust cost for aho_corasick should be dependent on the cost of its dependencies in some cases. One way to model this is with assumptions such as the following one: “if a crate’s dependencies are all safe to use, and the crate does not have any code flagged by a static analysis tool, then the crate is safe.” Cargo Sherlock uses a similar assumption to allow arguments that compose trust costs on a crate’s dependencies to obtain a trust cost for the crate itself.
2.5. Incorporating known risks
Finally, Ferris discovers the RustSec Advisory Database (Rust Secure Code Working Group, 2025), which lists vulnerabilities for different Rust crates. Ferris wants to check whether aho_corasick is present in the database. Unfortunately, nothing we have considered so far accounts for negative factors that would decrease trust in a crate; intuitively, this is because we have only incorporated assumptions that are about trustworthiness. We thus extend the tool with a distrust model to quantify known risks; we combine the results of the trust and distrust models to obtain a final label for the crate. In Figure 1, the distrust model is shown alongside the trust model with output in red.
3. Minimum Trust Problem
In this section, we present the formal model behind Cargo Sherlock, which aims to satisfy the requirements from Sections 1 and 2. We define the minimum trust problem formally: given a set of candidate assumptions and a conclusion, it asks to find the set of assumptions with minimum cost that entails the conclusion. We begin with an example, followed by two algorithms for the problem: first, a naive algorithm (Algorithm 1) and second, one based on Horn clause inference (Algorithm 2) that forms the basis for our trust solver implementation. To conclude the section, we show that the complexity of the minimum trust problem for Horn clauses is NP-complete, justifying such SMT-based approaches.
3.1. Problem definition
Propositional logic basics.
A propositional variable represents a true or false statement that is indivisible into other true or false statements, such as “the library is safe.” We use lowercase letters to denote propositional variables. A formula in propositional logic is either a propositional variable, constant ($\top$ for true, $\bot$ for false), or built from propositional variables and logical connectives (“and” $\wedge$, “or” $\vee$, “not” $\neg$, “implies” $\implies$). We use uppercase letters to denote formulas and uppercase Greek letters to denote sets of formulas. A Horn clause is a formula of the form $(p_{1}\wedge\dots\wedge p_{n})\implies q$, $\top\implies q$, or $(p_{1}\wedge\dots\wedge p_{n})\implies\bot$. The antecedent of the conditional is called the body and the consequent is called the head. A Horn clause is a definite clause if it is in the form $(p_{1}\wedge\dots\wedge p_{n})\implies q$. A set of definite Horn clauses is acyclic if there exists a function $f$ from propositional variables to natural numbers such that for every definite Horn clause $(p_{1}\wedge\dots\wedge p_{n})\implies q$, $f(q)>f(p_{i})$ for all $i$. A truth assignment is an assignment of true or false to propositional variables. A set of formulas $\Gamma$ entails another formula $A$ (written as $\Gamma\models A$) if and only if there is no truth assignment that makes all formulas in $\Gamma$ true and $A$ false. Equivalently, $\Gamma\models A$ if and only if $\Gamma\cup\{\neg A\}\models\bot$. A set of formulas $\Gamma$ is satisfiable if there exists some truth assignment that makes all formulas in $\Gamma$ true.
Definition 0 (Minimum trust problem).
Given as input (1) a set of propositional formulas $\Gamma=\{A_{1},\ldots,A_{n}\}$ representing assumptions, (2) a propositional variable $c$ representing a conclusion, and (3) a function $C\colon\Gamma\to\mathbb{N}$ that assigns a cost to making each assumption, output a set of assumptions $\Delta\subseteq\Gamma$ such that the sum of the costs of assumptions in $\Delta$ is minimized, subject to the constraint $\Delta\models c$.
We provide a simple example of the minimum trust problem, continuing the example from Section 2. Suppose that aho_corasick depends on memchr, where aho_corasick was developed by Alice, and memchr was developed by Bob and has no dependencies. Table 1 gives a set of candidate assumptions $A_{1},\ldots,A_{8}$ a user can make about the safety of the aho_corasick library and their associated costs in this scenario. We are asked to find a subset of $A_{i}$ with minimum cost which entails $c$. In this example, the minimum cost is $45$, incurred from assumptions $a$ (cost 20), $b$ (cost $5$), $(a\land m)\implies c$ (cost 10), and $b\implies m$ (cost $10$). Note that the set of assumptions is not necessarily unique in general (as there may be a tie for the minimum cost), though it is unique in this example.
No. Assumption Encoding ($A_{i}$) Cost
1 aho_corasick is safe. $c$ 100
2 memchr is safe. $m$ 100
3 aho_corasick has many downloads. $d$ 30
4 If aho_corasick has many downloads, then it is safe. $d\implies c$ 20
5 Alice is trustworthy. $a$ 20
6 Bob is trustworthy. $b$ 5
7 If Alice is trustworthy and memchr is safe, then aho_corasick is safe. $(a\land m)\implies c$ 10
8 If Bob is trustworthy, then memchr is safe. $b\implies m$ 10
Table 1. Example instance of the minimum trust problem.
Candidate assumptions, their encodings, and costs for the running example.
3.2. Naive algorithm
Algorithm 1 shows a naive solution for the minimum trust problem using an SMT solver by encoding the constraints as sentences over linear integer arithmetic and using quantifiers over propositional variables. To associate each assumption being made to a cost, we use auxiliary propositional variables $a_{1},\ldots,a_{n}$ to track each assumption $A_{1},\ldots,A_{n}$. Each $a_{i}$ is assigned to true if and only if $A_{i}$ is included in the solution set of assumptions $\Delta$. In the running example, we would have $a_{1}\implies c$, $a_{4}\implies(d\implies c)$, etc. We then encode the cost of the assumptions in $\Delta$ as a sum of if-then-else (iteite) terms; in the running example, the cost of any particular solution is given by $\text{{{cost}}}(a_{1},\ldots,a_{8})=ite(a_{1},100,0)+\dots+ite(a_{8},10,0)$. The constraint that the conclusion $c$ follows from the assumptions being made, $\Delta\models c$, is encoded as a formula $\text{{{valid}}}(a_{1},\ldots,a_{n})$. In the running example, $\text{{{valid}}}(a_{1},\ldots,a_{8})=\neg\exists c,m,d,a,b\>((a_{1}\implies c)\wedge\dots\wedge(a_{8}\implies(b\implies m))\wedge\neg c)$. Finally, we encode using quantifiers the fact that $\text{{{cost}}}(a_{1},\ldots,a_{n})$ is minimized subject to the constraint $\text{{{valid}}}(a_{1},\ldots,a_{n})$. In the running example, the SMT solver would produce a model that assigns $a_{5},a_{6},a_{7},a_{8}$ to true and $a_{1},a_{2},a_{3},a_{4}$ to false, incurring a cost of $45$.
Algorithm 1 Naive minimum trust algorithm.
1:  library: a software dependency in a package repository (e.g., a Rust crate on crates.io)
2:  get_all_assumptions(library): a function that returns a set of all possible assumptions for library
3:  $C$: a function that assigns a cost to an assumption
4:  S: an SMT solver over the theory of linear integer arithmetic
5:  procedure JoinAssumptions(lib)
6:    $F\leftarrow\top$
7:    $\Gamma\leftarrow\text{{{get\_all\_assumptions(lib)}}}$
8:    for all $A_{i}$ in $\Gamma$ do
9:     $F\leftarrow F\wedge(a_{i_{\text{{{lib}}}}}\implies A_{i})$
10:    for all dep in lib.dependencies do
11:     $F\leftarrow F\wedge\text{{{JoinAssumptions}}}(\text{{{dep}}})$ $\triangleright$ JoinAssumptions is called recursively on each dependency
12:    output $F$
13:  procedure SolveMinimumTrust()
14:    $F(a_{1},\ldots,a_{n})\leftarrow\text{{{JoinAssumptions(library)}}}$
15:    $\{v_{1},\ldots,v_{m}\}\leftarrow\text{{{atoms}}}(F)\setminus\{a_{1},\ldots,a_{n}\}$ $\triangleright$ atoms returns the set of propositional variables in a formula
16:    $\text{{{valid}}}(a_{1},\ldots,a_{n})\leftarrow\neg\exists c\exists v_{1}\ldots\exists v_{m}\>(F(a_{1},\ldots,a_{n})\wedge\neg c)$ $\triangleright$ $c$ is the conclusion ‘the library is safe’
17:    $\text{{{cost}}}(a_{1},\ldots,a_{n})\leftarrow\sum_{i=1}^{n}ite(a_{i},C(a_{i}),0)$
18:    $\text{{{minimal}}}\leftarrow\forall a_{1}\ldots\forall a_{n}\>(\text{{{valid}}}(a_{1},\ldots,a_{n})\implies\text{{{cost}}}(a^{*}_{1},\ldots,a^{*}_{n})\leq\text{{{cost}}}(a_{1},\ldots,a_{n}))$
19:    $\text{{{S.solve}}}(\text{{{valid}}}(a^{*}_{1},\ldots,a^{*}_{n})\land\text{{{minimal}}})$ $\triangleright$ $(a^{*}_{1},\ldots,a^{*}_{n})$ is the solution to the minimum trust problem
[…]
Theorem 2.
Algorithm 1 is correct.
Proof.
Algorithm 1 recursively calls JoinAssumptions on each library in the dependency tree on line 11 to construct the formula $F(a_{1},\ldots,a_{n})$ that represents the assumptions being made, then uses the above encoding to create the constraint $\text{{{valid}}}(a_{1},\ldots,a_{n})$. The algorithm encodes that the cost of assumptions is minimized subject to the constraint $\text{{{valid}}}(a_{1},\ldots,a_{n})$ on line 18. ∎
Although it is correct, the use of quantifiers makes the first algorithm computationally expensive; it generates a linear number of nested quantifiers in the size of the dependency tree, which have to be expanded before applying a SAT solver. While some of this complexity is unavoidable (see Section 3.4), to address this we next propose an algorithm that does not use quantifiers, and which we have found to be more effective in practice.
3.3. Horn clause algorithm
The key idea of Algorithm 2 is to restrict the set of assumptions to be acyclic definite Horn clauses; in this case, the problem can be solved efficiently with a binary search reduction using quantifier-free queries to an SMT solver. The set of assumptions that can be expressed as acyclic definite Horn clauses is sufficient for all of the assumptions we identify based on prior work (Table 2(a)).
In the running example, observe that the set of conjuncts of $F$ that establish the safety of the aho_corasick library ($c$) are $a_{1}\implies c$, $a_{4}\implies(d\implies c)$, and $a_{7}\implies((a\land m)\implies c)$. Rewritten as Horn clauses with a body and a head, these conjuncts are $a_{1}\implies c$, $(a_{4}\land d)\implies c$, and $(a_{7}\land a\land m)\implies c$. Therefore, with the goal of showing that $c$ is true, the formula $a_{1}\lor(a_{4}\land d)\lor(a_{7}\land a\land m)$ must evaluate to true. The only conjunct of $F$ that establishes $d$ is $a_{3}\implies d$. Therefore, we can substitute $d$ with $a_{3}$ in the formula, resulting in $a_{1}\lor(a_{4}\land a_{3})\lor(a_{7}\land a\land m)$. By repeating this unfolding process, we can simplify the formula to $a_{1}\lor(a_{4}\land a_{3})\lor(a_{7}\land a_{5}\land a_{6}\land a_{8})$. This formula will evaluate to true if and only if the assumptions made entail the conclusion $c$. Algorithm 2 uses this observation to solve the minimum trust problem when the assumptions are acyclic definite Horn clauses. Due to binary search, Algorithm 2 produces a logarithmic number of calls to the SMT solver with respect to the maximum cost, and each call is quantifier-free.
Algorithm 2 Horn clause optimized minimum trust algorithm.
1:  library: a library
2:  $C$: a function that assigns a cost to an assumption
3:  S: an SMT solver over the theory of linear integer arithmetic
4:  procedure Unfold($F$)
5:    $F\leftarrow\text{{{simplify}}}(F)$ $\triangleright$ simplify the formula into a conjunction of Horn clauses
6:    $\Pi\leftarrow\text{{{clauses}}}(F)$ $\triangleright$ get the set of Horn clauses in $F$
7:    $\leftarrow\bot \triangleright EE$ is to be constructed such that $E$ is true iff the assumptions entail the conclusion
8:    for all $\pi_{i}$ in $\Pi$ do
9:     if $\text{{{head}}}(\pi_{i})$ is the conclusion $c$ then
10:       $E\leftarrow E\lor\text{{{body}}}(\pi_{i})$ $\triangleright$ collect all the bodies of the Horn clauses with head $c$
11:    $\leftarrow\text{{{atoms}}}(F)\setminus\{a_{1},\ldots,a_{n}\} \triangleright VV$ stores the set of propositional variables in $F$ other than the $a_{i}$s
12:    for all $v_{i}$ in $V$ do
13:     $T\leftarrow\bot$ $\triangleright$ the formula to be substituted in place of $v_{i}$
14:     for all $\pi_{i}$ in $\Pi$ do
15:       if $\text{{{head}}}(\pi_{i})$ is $v_{i}$ then
16:        $T\leftarrow T\lor\text{{{body}}}(\pi_{i})$ $\triangleright$ collect all the bodies of the Horn clauses with head $v_{i}$
17:       $E\leftarrow E[v_{i}/T]$ $\triangleright$ substitute $T$ in place of $v_{i}$ in $E$
18:     for all $\pi_{i}$ in $\Pi$ do
19:       $\pi_{i}\leftarrow\pi_{i}[v_{i}/T]$ $\triangleright$ substitute $T$ in place of $v_{i}$ in $\pi_{i}$
20:    output $E$
21:  procedure SolveHornMinimumTrust()
22:    $F(a_{1},\ldots,a_{n})\leftarrow\text{{{JoinAssumptions(library)}}}$ $\triangleright$ same as in Algorithm 1
23:    $\text{{{valid}}}(a_{1},\ldots,a_{n})\leftarrow\text{{{Unfold}}}(F)$
24:    $\text{{{cost}}}(a_{1},\ldots,a_{n})\leftarrow\sum_{i=1}^{n}ite(a_{i},C(a_{i}),0)$
25:    $\text{{{left}}}\leftarrow 0$ $\triangleright$ initialize the min cost as 0
26:    $\text{{{right}}}\leftarrow\sum_{i=1}^{n}C(a_{i})$ $\triangleright$ the max cost is the sum of the costs of all assumptions
27:    while $\text{{{left}}}<\text{{{right}}}$ do
28:     $k\leftarrow\lfloor(\text{{{left}}}+\text{{{right}}}+1)/2\rfloor$ $\triangleright$ binary search for min cost
29:     $\text{{{cost\_leq\_k}}}\leftarrow\text{{{S.solve}}}(\text{{{valid}}}(a_{1},\ldots,a_{n})\wedge\text{{{cost}}}(a_{1},\ldots,a_{n})\leq k)$
30:     if cost_leq_k is SAT then
31:       $\text{{{left}}}\leftarrow k$
32:     else
33:       $\text{{{right}}}\leftarrow k-1$
34:    $\text{{{min\_cost}}}\leftarrow\text{{{left}}}$
35:    output S.model[$a_{1},\ldots,a_{n}$], min_cost
Theorem 3.
Algorithm 2 is correct for assumptions consisting of acyclic definite Horn clauses.
Proof.
Algorithm 2 first constructs $E$ as the disjunction of all the bodies of the Horn clauses whose heads are the conclusion $c$ (line 10). The conclusion is provable from the assumptions if and only if $E$ evaluates to true. Algorithm 2 then proceeds to iterate through $V$ with the goal of eliminating all the propositional variables in $V$ from $E$. For each propositional variable $v_{i}$ in $V$, the algorithm constructs a formula $T$ consisting of the disjunction of the bodies of the Horn clauses whose heads are $v_{i}$ and then eliminates $v_{i}$ by substituting $T$ in place of $v_{i}$ in $E$, similar to the construction of $E$ on line 10. After the substitution, $E$ will still be a necessary and sufficient condition for the conclusion to be true. The substitution is then applied to each remaining clause in line 19; since the Horn clauses are acyclic, eliminating $v_{i}$ is guaranteed to remove $v_{i}$ from all future substitutions. This process is repeated until all propositional variables in $V$ are eliminated from $E$. Each step in Unfold maintains that $E$ is necessary and sufficient for the conclusion to be true. The formula generated by Unfold is used in a binary search to repeatedly query S for valid minimum trust solutions with different costs until the minimum cost is found. ∎
It should be noted that the unfolding procedure in Algorithm 2 is similar to the well-known SLD resolution algorithm for logic programming (Kowalski, 1974). Unfortunately, the generated formula has an exponential number of clauses in the worst case (with respect to the number of crates in the dependency tree). This formula, however, is in disjunctive normal form (DNF) after minor simplification, which means the SMT solver can solve each clause independently. Note that it is also possible to have a reduction that produces a polynomial-size formula, though this is not the reduction we use in our implementation.
3.4. Complexity
To establish the complexity of the minimum trust problem (Definition 1), we study the corresponding decision problem. To that end, define MINTRUST to be the problem that asks, given a tuple $(\Gamma,c,C,k)$ — consisting of a set of assumptions $\Gamma$, a conclusion $c$, a cost function $C$, and integer $k$ — does there exists a set of assumptions $\Delta\subseteq\Gamma$ such that the sum of the costs of the assumptions in $\Delta$ is less than or equal to $k$ and $\Delta\models c$? Similarly, define HORN-MINTRUST to be the restriction of this problem where $\Gamma$ consists of Horn clauses. The following theorems establish that the unrestricted minimum trust problem is computationally infeasible (Theorem 4) and that the Horn-restricted minimum trust problem is intractable but still within the reach of SMT solvers (Theorem 5).
Theorem 4.
MINTRUST is $\Sigma_{2}^{P}$-hard.
Proof.
We reduce from Shortest Implicant Core (SIC), which is known to be $\Sigma_{2}^{P}$-hard (Schaefer and Umans, 2002) (Umans, 2001). The SIC problem asks: given a DNF formula $\varphi$, an implicant $I$ of $\varphi$ (a set of literals whose conjunction implies $\varphi$), and an integer $k$, does an implicant $I^{\prime}\subseteq I$ of size at most $k$ exist? Given an instance of SIC with a DNF formula $\varphi$, an implicant $I$ of $\varphi$, and an integer $k$, we construct a MINTRUST instance $(\Gamma,c,C,k)$. Let $\Gamma$ contain an assumption for each literal in $I$ with a cost of $1$ and the assumption $\varphi\implies c$ with a cost of 0.
By construction, $\Delta\models c$ if and only if $\Delta\models\varphi$ and $(\varphi\implies c)\in\Delta$. Furthermore, $\Delta\models\varphi$ if and only if $\Delta$ contains an implicant $I^{\prime}\subseteq I$ of $\varphi$. Then, the cost associated with $\Delta$ is equal to the size of $I^{\prime}$, so $\varphi$ has an implicant of size at most $k$ if and only if the MINTRUST instance has a solution of cost at most $k$. Therefore, an implicant $I^{\prime}\subseteq I$ of $\varphi$ of size at most $k$ exists if and only if this MINTRUST instance has a solution of cost at most $k$. ∎
Theorem 5.
HORN-MINTRUST is NP-complete.
Proof.
HORN-MINTRUST is NP-hard. We reduce from Vertex Cover (Karp, 1972). Given an instance of vertex cover with a graph $G=(V,E)$ and an integer $k$, we construct a HORN-MINTRUST instance $(\Gamma,c,C,k)$. Let $\Gamma$ contain an assumption $v_{i}$ for each vertex $v_{i}\in V$ with cost $1$, an assumption $v_{i}\implies e_{j}$ for each vertex $v_{i}\in V$ that is incident to edge $e_{j}\in E$ with cost 0, and an assumption $(e_{1}\land\dots\land e_{|E|})\implies c$ with cost 0. By construction, $\Delta\models c$ if and only if $\Delta\models e_{j}$ for all $e_{j}\in E$ and $((e_{1}\land\dots\land e_{|E|})\implies c)\in\Delta$. Furthermore, $\Delta\models e_{j}$ for each $e_{j}\in E$ if and only if $v_{i}\in\Delta$ and $(v_{i}\implies e_{j})\in\Delta$ for some incident $v_{i}$. It follows that $\Delta\models e_{j}$ for all $e_{j}\in E$ if and only if $\Delta$ is a subset of the $v_{i}$s for a vertex cover of $G$, together with the assumptions for incident edges. The cost associated with $\Delta$ is equal to the size of the vertex cover, which is at most $k$.
HORN-MINTRUST is in NP. We construct a verifier for HORN-MINTRUST. On input a candidate solution $\Delta$, the verifier checks that the sum of the costs of the assumptions in $\Delta$ is at most $k$ in one pass. Since $\Delta\models c$ holds if and only if the conjunction of the Horn clauses in $\Delta$ with $(c\implies\bot)$ is unsatisfiable, and Horn-satisfiability is in P, checking that $\Delta\models c$ is possible in polynomial time. ∎
4. Design and Implementation
Cargo Sherlock is a Python-based tool for quantifying trust in the Rust supply chain ecosystem, given as input a Rust crate. [^3] https://github.com/davispl/cargo-sherlock It provides a logs mode, which collects and displays information about the crate, and a trust mode (see output in Figure 1), which is used to compute the trust cost and severity label of the crate, with components as detailed in the following subsections.
Figure 2. High level overview of Cargo Sherlock.
High level overview of the Cargo Sherlock tool. The tool gathers information from multiple sources, builds a relational model, and uses Z3 solver to compute the trust cost and severity label of a given Rust crate.
4.1. Data collection
Cargo Sherlock first gathers crate metadata from multiple sources. From crates.io, it collects information about the author(s), download counts, dependencies, and GitHub repository stars and forks. From Cargo Vet (Mozilla, 2024), it sources audit reports from organizations including Google, Mozilla, and Bytecode Alliance. Cargo Sherlock also gathers the outputs from static analysis tools (Miri (Rust team, 2024) and Cargo Scan (PL and PLSysSec, 2026; Zoghbi et al., 2026)) and vulnerability information from RustSec (Rust Secure Code Working Group, 2025).
4.2. Relational model
The aggregated data is organized into a relational model as shown in Figure 3. The primary entities include: Crates, Authors, Audits, Dependencies, and Tool Results. Each Crate includes basic metadata such as its name, version, downloads, and related git data (e.g., stars and forks). Crates can have multiple associated tool results, audits, authors, and dependencies, following a one-to-many pattern. Dependencies connect crates to other crates, while Audits link organizations and trust criteria to crates.
Refer to caption Entity-Relationship (ER) diagram of the Cargo Sherlock relational model, showing entities such as Crate, Author, Audit, Dependency, and Tool Results, along with their attributes and relationships.
Figure 3. Cargo Sherlock entity-relationship (ER) model.
4.3. Selection of assumptions
Our candidate assumptions are motivated by prior work on software supply chain trust. Hamer et al. (Hamer et al., 2025) identified key trust factors among Rust developers, while Boughton et al. (Boughton et al., 2024) decomposed these into distinct dimensions. Table 2 shows some of the assumptions we support and their justifications in prior research.
Cargo Sherlock supports three kinds of assumptions: positive, negative, and parameterized. The costs assigned to each candidate assumption are configurable by the user. In our implementation, we provide default costs, which we arrived at through a process of trial and error.
(a) Trust model assumptions
Cost Assumption Justification
100 The crate is safe. Base
25–100 If the crate has many downloads, it is safe. (Hamer et al., 2025) (Tbl. 7)
5 If the crate version has a passed audit, it is safe. (Hamer et al., 2025) (Tbl. 7)
20 If a past version has a passed audit, it is safe. (Hamer et al., 2025) (Tbl. 7)
20–100 If the repository has many stars/forks, it is safe. (Hamer et al., 2025) (Tbl. 7)
10 If the crate has no side effects and all dependencies are safe, it is safe. (Terei et al., 2012; Zoghbi et al., 2026)
5 If the author of the crate is trusted, it is safe. (Hamer et al., 2025) (Tbl. 7)
(b) Distrust model assumptions
Cost Assumption Justification
100 The crate is not safe. Base
5 If the crate appears on RustSec with a critical severity label, it is not safe. (Boughton et al., 2024)
90 If the crate appears on RustSec and the provided crate version is patched, it is not safe. (Boughton et al., 2024)
30 If the crate is flagged by Miri, it is not safe. (Boughton et al., 2024)
60–100 If the crate has many side effects, it is not safe. (Terei et al., 2012; Zoghbi et al., 2026)
10 If the crate has an unsafe dependency, it is not safe. (Boughton et al., 2024)
Table 2. Selected assumptions used by Cargo Sherlock (including justifications in prior work).
Side-by-side tables summarizing trust and distrust assumptions with costs and justifications.
Positive assumptions:
These are candidate assumptions a user might use to reason that a crate is safe. By default, Cargo Sherlock includes the following:
  • Trusted Author: A crate is safe if it is authored by a well-established developer in the Rust community (Hamer et al., 2025). Cargo Sherlock includes a default list of trusted developers, which users can customize.
  • Audit Passed: A crate is safe if it has passed an audit (Hamer et al., 2025). This assumption has a variant for crates with a previous version that have passed an audit.
  • Side Effects and Safe Dependencies: A crate is safe if it has no side effects, i.e., no observable change to program state or the external environment beyond returning a value, and all of its dependencies are safe (Terei et al., 2012; Zoghbi et al., 2026). Side effects are identified using Cargo Scan (PL and PLSysSec, 2026).
  • Crate is Safe: By default, a crate is safe (this assumption has the highest possible cost). Cargo Sherlock includes this assumption with no justification as a base case.
Negative assumptions:
These are candidate assumptions a user might use to reason that a crate is unsafe. By default, Cargo Sherlock includes the following:
  • RustSec: A crate is unsafe if it appears in the RustSec advisory database (Rust Secure Code Working Group, 2025). This assumption has variants for crates with a previous version that appear in RustSec and for different RustSec labels.
  • Tool Results: A crate is unsafe if it is flagged by a static analysis tool (Boughton et al., 2024). Currently, Cargo Sherlock supports the assumption that crates flagged by Miri (Rust team, 2024) are unsafe.
  • Unsafe Dependency: A crate is unsafe if it has an unsafe dependency. (Boughton et al., 2024)
  • Crate is Unsafe: By default, a crate is unsafe (this assumption has the highest possible cost). Cargo Sherlock includes this assumption with no justification as a base case.
Parameterized assumptions:
These are candidate assumptions for which the trust cost is calculated as a function of certain metadata parameters; we use a smoothing function so that the trust cost changes gradually with changes to the parameter. We also allow parameterized negative assumptions, which work for distrust cost in the same way. By default, Cargo Sherlock includes the following:
  • Number of Downloads: A crate with many downloads is safe (Hamer et al., 2025).
  • GitHub Stars and Forks: A crate with many stars and forks on its GitHub is safe (Hamer et al., 2025).
  • Side Effects: A crate with many side effects is unsafe (Terei et al., 2012; Zoghbi et al., 2026).
Cargo Sherlock applies a consistency checker to ensure checks on all assumption costs; currently, it checks that no assumption has a cost higher than the base cost of 100.
4.4. Computing the trust cost
We use the Z3 SMT solver (de Moura and Bjørner, 2008) to compute minimum trust and distrust costs using either Algorithm 1 or Algorithm 2 (default). In our implementation, we replace the explicit binary search over costs with Z3’s optimizing SAT interface to obtain the minimum cost directly. The trust cost ranges from 0 to 100, where a lower trust cost represents a crate that is easier to trust. Similarly, the distrust cost ranges from 0 to 100, where a lower distrust cost represents a crate that is easier to distrust. The final label is derived by combining both costs, using the function illustrated in Figure 4.
Figure 4. Cargo Sherlock combines trust and distrust costs to produce a severity label.
A scatter plot showing the relationship between distrust cost and trust cost, divided into regions labeled Safe, Low Severity, Medium Severity, High Severity, and Critical. The plot illustrates how Cargo Sherlock combines these costs to produce a severity label for Rust crates.
5. Evaluation
In this section, we evaluate Cargo Sherlock on the following research questions:
  • RQ1: Does Cargo Sherlock detect synthetically generated typosquatting attacks?
  • RQ2: Does Cargo Sherlock detect real-world supply chain risks?
  • RQ3: How does Cargo Sherlock output fare on known vulnerable crates?
  • RQ4: How does the performance of Algorithms 1 and 2 scale with the number of dependencies?
Experimental setup.
All evaluations were conducted on an M2-Pro MacBook Pro with 16 GB of RAM. Our experiments are run on various sets of crates as described below: the 100 most downloaded crates from crates.io, a random sample of 1000 crates, a sample of crates flagged in the RustSec database, and two case studies. All experiments were run between March and November 2025.
User model.
For all experiments, we adopt an idealized “rational developer” model in the sense of Dimoulas et al. (Dimoulas and Felleisen, 2025). The rational developer interacts with our tool as follows: when they intend to add a dependency to their existing Rust project, they execute Cargo Sherlock on the selected crate. This generates a report accompanied by a severity label; we assume that the developer looks at the report and decides whether to go forward with adding the dependency based on the severity label. For each experimental section, we interpret our findings through this user model and discuss the resulting implications for prospective users of Cargo Sherlock.
5.1. RQ1: Does Cargo Sherlock detect synthetically generated typosquatting attacks?
Typosquatting is an attack technique that exploits typographical errors to deceive users into downloading a malicious package, such as downloading serde_yml in place of serde_yaml. We consider typosquatting because it is a common category of supply chain attack, and aligns closely with Cargo Sherlock’s intended use case. To address RQ1, we constructed synthetic typosquatted versions of the top 100 most frequently downloaded Rust crates from crates.io. We selected popular crates as a test sample because their popularity makes them a lucrative target for an attacker. We generated typosquatted variants by modifying the name and author of each crate, while keeping the functionality unchanged. To determine how many downloads would be obtained by a lesser-known crate by default, we measured that placeholder crates on crates.io received an average of 171 downloads within 3 months. We use this number for the download count in our synthetic experiment. After creating typosquatted variants, we evaluated them using Cargo Sherlock and compared their severity labels with corresponding original crates. Figure 5(a) shows the distribution of obtained severity labels. Most of the original crates are classified as safe (73%) and no crate is classified higher than low_severity. On the other hand, among typosquatted crates, most are classified as critical. However, 25 typosquatted crates also receive the safe label, likely because they have no side effects and they all have safe dependencies. The severity label increases for the typosquatted crate when compared to the original crate in all other cases.
RQ1 Findings: Most typosquatted versions of popular Rust crates incur substantially higher trust costs and higher severity labels compared to their original counterparts when evaluated by Cargo Sherlock. Implication: A user interacting with the tool would identify the typosquatted crate variant when adding a dependency to their code.
(a) Cargo Sherlock results on synthetically generated typosquatting attack crates (compared to their real versions).
Cargo Sherlock severity labels on synthetically generated typosquat crates (compared to their real versions).
Crate Trust Cost Distrust Cost Severity Label
Pair 1: fast_log (original) vs faster_log (typosquatted)
fast_log-1.7.7 37 73 MEDIUM_SEVERITY
faster_log-1.7.8 75 80 CRITICAL
Pair 2: serde_yaml (original) vs serde_yml (AI-maintained fork)
serde_yaml-0.9.33 6 70 SAFE
serde_yml-0.0.12 32 70 LOW_SEVERITY
(b) Cargo Sherlock results on real-world incidents.
A table summarizing trust cost, distrust cost, and severity level for two pairs of crates: fast_log vs faster_log, and serde_yaml vs serde_yml.
Figure 5. Sections 5.1 and 5.2 experimental results.
RQ1 and RQ2 severity results overview.
5.2. RQ2: Does Cargo Sherlock detect real-world supply chain risks?
We address this question as a case study on Rust crates that have been involved in known incidents on crates.io. We obtained the source code for two such crates that exist to our knowledge: faster_log and serde_yml. [^4] We are aware of known incidents for at least two other Rust crates: rustdecimal (Rust Secure Code Working Group, 2024b) and asyncprintln (Pearce, 2025). Unfortunately, we were unable to obtain the source code for the typosquatted versions of these crates to include them in the analysis.
We first examine faster_log, an identified malicious typosquat of fast_log (Pearce, 2025). Beyond reproducing the original crate’s logging functionality, it injects code that scans log files for sensitive material (including cryptocurrency private keys) and exfiltrates any matches to a remote endpoint. Cargo Sherlock assigns a MEDIUM_SEVERITY label (trust cost $37$, distrust cost $73$) to fast_log, basing the trust on the author (they have a history of publishing crates) and the distrust on the presence of side effects (it performs file I/O). In contrast, Cargo Sherlock assigns faster_log a CRITICAL label, with a substantially higher trust cost ($75$) and elevated distrust cost ($80$). The change in authorship (the author had no other crates) and side effects leads to a different trust and distrust cost.
Our second case is serde_yml, a fork of serde_yaml (the latter authored by David Tolnay, a well-established and trusted developer in the Rust community). After the original crate was marked unmaintained, the fork was published on crates.io and had accumulated over 1,057,000 downloads by March 19, 2025 (Tolnay, 2025). The fork incorporates a substantial amount of AI-generated code, including low-quality “AI slop” changes that were later criticized by Tolnay (Tolnay, 2025). As shown in Table 5(b), the original crate receives a SAFE label, while the fork is classified with the higher LOW_SEVERITY, primarily due to the change in authorship. The distrust costs are similar across the two crates since no identifiable malicious code or side effects were introduced.
RQ2 Findings: Cargo Sherlock assigns higher severity labels to the malicious typosquat faster_log and AI-maintained fork serde_yml, relative to their original crates. Implication: Users interacting with Cargo Sherlock may perceive these crates as risky.
5.3. RQ3: How does Cargo Sherlock output fare on known vulnerable crates?
To address RQ3, we compare Cargo Sherlock’s severity labels with those from RustSec (Rust Secure Code Working Group, 2025), a well-known database of security advisories for Rust crates. RustSec classifies crates by severity levels: critical, high, medium, low, and informational. We analyzed 592 unique crate-version pairs listed in RustSec. [^5] This number is as of March 25, 2025. For this experiment, all negative assumptions related to RustSec were removed to determine what labels would have been assigned to vulnerable crates prior to them being known to be vulnerable.
Figure 6(a) illustrates the distribution of severity labels for crates listed on RustSec. The x-axis shows the severity labels assigned by RustSec, and each bar is split according to the labels assigned by Cargo Sherlock. While some crates receive a higher severity label, most crates are classified as SAFE or LOW_SEVERITY. For example, these include popular crates (hence, more likely to be marked SAFE) that are listed in RustSec with an INFO label due to being marked as unmaintained. Figure 6(b) confirms that this behavior is generally driven by the popularity of RustSec-listed crates: approximately 75% of RustSec crates belong to the top 10% of crates.io by downloads, and only around 15% of RustSec crates are found in the bottom 80%.
(a) Distribution of Cargo Sherlock vs. RustSec severity labels.
A bar chart showing the distribution of Cargo Sherlock versus RustSec severity labels on the crates listed in RustSec.
(b) Distribution of crates in RustSec by downloads.
A bar chart showing the percentile distribution of crates in RustSec by downloads on crates.io.
Figure 6. Section 5.3 experimental results.
RQ3 results plots.
RQ3 Findings: Crates that are marked SAFE by Cargo Sherlock are not necessarily free from vulnerabilities, especially in cases with a high number of downloads. Implication: Users interacting with Cargo Sherlock might perceive some vulnerable crates as safe.
5.4. RQ4: How does the performance of Algorithms 1 and 2 scale with the number of dependencies?
To address RQ4, we sampled 1,000 random crates from crates.io. We excluded 47 crates from our evaluation because all of their versions had been yanked, leaving 953 crates. To isolate algorithmic running time, we excluded data collection time from all measurements. The cumulative distribution function (CDF) plot in Figure 7(a) shows the runtime for both algorithms, using a timeout of 600 seconds. Algorithm 1 successfully evaluates 298 crates and timed out on 500 crates. It crashes on 155 crates, out of which 150 crashes are due to the system’s maximum recursion depth being exceeded. The remaining 5 crashes are shared by both algorithms and are due to parsing failures. Algorithm 2 successfully evaluated 900 crates and only timed out on 48 crates.
(a) CDF of running time for Algorithm 1 vs. Algorithm 2.
[…]
Figure 7. Section 5.4 scalability results.
RQ4 results plots.
The scatter plot (Figure 7(b)) shows that Algorithm 1 rapidly becomes infeasible as the number of dependencies increases, with most evaluations timing out beyond approximately 10 dependencies. Conversely, Algorithm 2 scales better, successfully evaluating crates with up to 150 dependencies. A regression fit suggests that the scaling of the Horn clause algorithm is exponential. This is expected, as unfold in Algorithm 2 generates an exponential DNF formula in the worst case.
RQ4 Findings: Algorithm 1 becomes infeasible after 10 dependencies, while Algorithm 2 scales to at least 150 dependencies. Implication: Users can run Cargo Sherlock on crates with a large number of dependencies and receive results within 10 minutes.
6. Related Work
Supply chain security:
Software supply chain attacks (Ohm et al., 2020; Wang, 2021; Ohm and Stuke, 2023) such as the SolarWinds attack (Martínez and Durán, 2021) and XZ Utils backdoor (Zorz, 2024; Freund, 2024) have received attention from researchers and industry. Tools like Crev (development team, 2024b) and JFrog (Chin, 2023) provide support for auditing dependencies and tracking associated metadata. Software bill of materials (SBOM) (Muirí, 2019; Hendrick and Zemlin, 2022; Xia et al., 2023; Zahan et al., 2023) is a general set of techniques for providing an overview of software contents, which can be used for human understanding; these are similar to the relational model used in our work. Tools have also been built for specific languages such as JavaScript (Zimmermann et al., 2019; Cogo et al., 2019; Zahan et al., 2022) and Go (Google, 2023b, a). Especially relevant to our work, a few techniques in this space have used formal methods: Koscinski and Mirakhorli (Koscinski and Mirakhorli, 2024) propose a technique to formally model Common Weakness Enumerations (CWEs) using an encoding in the Alloy specification language. Software configuration management (SCM) (Zeller and Snelting, 1997) is also related; it aims to identify a configuration of dependencies that satisfies formal constraints. Our work can be thought of as a similar configuration problem to optimize for trust cost, and selecting among multiple possible configurations (different versions of different software packages) would be an interesting direction for future work.
Safety and security for Rust code:
RustSec (Rust Secure Code Working Group, 2025) and Lib.rs (Lesiński, 2024) provide helpful databases of known vulnerabilities and libraries. Cargo Vet (Mozilla, 2024), Cargo Audit (Rust Secure Code Working Group, 2024a), Cargo Crev (development team, 2024a), and Cargo Scan (PL and PLSysSec, 2026; Zoghbi et al., 2026) (which inspired the side effects metrics used in our tool) are tools to assist with human auditing for Rust code. Formal verification for Rust includes work on foundations — like RustBelt (Jung et al., 2017), Oxide (Weiss et al., 2019), BoCa (Wagner et al., 2025), Stacked/Tree Borrows (Jung et al., 2019; Villani et al., 2025), and others (Grannan et al., 2025) — and tools such as Verus (language team, 2022), Prusti (Astrauskas et al., 2019), Kani (VanHattum et al., 2022), and Flux (Lehmann et al., 2023). Program analysis tools like MIRChecker (Li et al., 2021), Miri (Rust team, 2024), and Rudra (Bae et al., 2021) detect memory safety and undefined behaviors in Rust code, and fuzzing tools such as SyRust (Takashima et al., 2021) and RULF (Jiang et al., 2022) use automated testing to identify bugs. Sandboxing tools (e.g., (Bang et al., 2023; Lamowski et al., 2017; Kirth et al., 2022; Liu et al., 2020; Almohri and Evans, 2018; Rivera et al., 2021; Gulmez et al., 2023)) typically aim to protect Rust code from its unsafe subset. However, even the safe subset of Rust can be exploited, and this is sometimes missed by prior tools (Hassnain and Stanford, 2024; user), 2024; Zoghbi et al., 2026), which motivates work incorporating human factors.
Trust in software engineering:
Hamer et al. (Hamer et al., 2025) explored the impact of contributor reputation and other indicator metrics in ecosystems like Rust and GitHub. Boughton et al. (Boughton et al., 2024) proposed a framework to decompose and quantify trust within open-source software. We used these studies to motivate the specific assumptions that we support in our minimum trust cost model. Devanbu et al. (Devanbu et al., 1998) argue for a combination of trusted hardware, key management, and verification techniques to elevate trust. Toone et al. (Toone et al., 2003) introduce trust mediation, a framework for managing complex distributed trust relationships. Outside of computer science, models of trust have been studied in other settings (e.g., (Mayer et al., 1995)).
Acknowledgements.
This research was supported in part by the US National Science Foundation under award CCF #2327338. The authors would like to thank the anonymous reviewers and Prem Devanbu for feedback leading to improvements to the paper, and Audrey Gobaco for helpful discussions and work on related projects.